OpenStudy (anonymous):

Find an equation that models the path of a satellite if its path is a hyperbola, a = 55,000 km, and c = 81,000 km. Assume that the center of the hyperbola is the origin and the transverse axis is horizontal.

OpenStudy (anonymous):

My work: a^2 + b^2 = c^2 55,000^2 + b^2 = 81,000^2 b = 59464.27 (x^2/a^2) - (y^2/b^2) = 1 x^2/55,000^2 - y^2/59464^2 = 1 x^2/3,025,000,000 - y^2/3,535,967,296 = 1